# Bike-Bench: A Bicycle Design Benchmark for Generative Models with Objectives and Constraints

### v1

@@ 5 Benchmarking Metrics and Procedure
The design objectives detailed above yield a set of 10 design objectives and 15 design constraints, which are explained in detail in the Appendix A.4. However, to compare the performance of different generative models, we desire a concise set of summary metrics to evaluate sample sets. This section describes Bike-Bench’s metrics and general benchmarking procedure.
@@ 5.1 Metrics
Though there are countless possibilities for generative model evaluation metrics [24], we select three: validity, optimality, and similarity, which we describe below. Unlike other benchmarks with singular objectives, models benchmarked on Bike-Bench should be compared using all three scoring metrics.
Validity (Constraint Satisfaction Rate):
To evaluate a model’s constraint satisfaction ability, we measure the proportion of generated designs that simultaneously satisfy all constraints.
Optimality (Hypervolume Metric):
We calculate the hypervolume metric over any designs that simultaneoulsy satisfy all constraints. The Hypervolume metric is a staple of multi-objective optimization literature which calculates the overall multi-objective optimality of a set of designs. Hypervolume generally benefits from larger and more diverse solution sets. More details on hypervolume metric and its calculation are included in the Appendix.
Similarity (Maximum Mean Discrepancy):
To capture a model’s ability to match the manifold of existing designs, we evaluate distributional similarity to a subset of the BIKED dataset withheld from the model during training. We select Maximum Mean Discrepancy, a common kernel-based statistical discrepancy measure used to compare sets of samples. Similarity remains an important facet of design synthesis since it helps enforce two goals: ‘realism’ and ‘design space coverage’ [24].
This medley of performance scores roughly capture a model’s ability to match the manifold of existing designs, avoid constraint violation and optimize design objectives. The function of validity and optimality are self-explanatory for design – maximally useful designs must be both feasible and optimized. The role that manifold similarity plays in design synthesis is less obvious. Just as human designers guide their decision-making based on the space of existing designs, design-space modeling can also assist algorithmic design synthesis approaches, a key motivation for the use of generative models. Statistical similarity serves to enforce two desirable attributes in generative models for design. The first is ‘design realism’ – what does it mean to be a bike and not something entirely different, say, a wheelbarrow? The second desirable attribute is ‘design space coverage, ’ encouraging models to learn more than just a small niche subspace of the overall breadth of bicycle designs, expanding their utility.
@@ 5.2 Benchmarking Procedure
Bike-bench’s principal benchmark evaluates conditional generative modeling, providing 10,000 conditional cases (text prompt, rider anthropometric dimensions, use case, and reference image) and collecting one sample per condition from the generative model. Distributional similarity, constraint satisfaction rate over all 15 constraints, and hypervolume metric over all 10 evaluation objectives are evaluated for the 10,000 resulting designs and their respective conditions.
We also provide an unconditional generative modeling benchmark, where generative models are trained to satisfy a singular fixed condition, then are queried for 1,000 samples. To make scores less sensitive to specific conditions, 10 unique conditional cases are given, models are trained 10 times, and mean scores over these 10 conditional cases are reported.
To avoid bias, models are trained only on the 3600 training samples from the original 4500 human-designed bicycles, and not on synthetically-generated designs (The synthetic dataset is used exclusively for evaluation of aesthetics scores). Models are allowed to train on either a mixed-datatype or continuous version of the data. Designs generated by models trained on continuous data are mapped back to a mixed-datatype representation prior to evaluation – booleans are decided using rounding, while one-hot categorical data is decided using argmax.
@@ 6 Models Benchmarked
Our principal goal in the benchmarking of baselines is to span a variety of types of design generation procedures ranging from LLMs to tabular generative models to optimization algorithms. Because we do not focus on extensively comparing methods within specific classes, we limit our benchmarking to just one or a few methods per class. The various features and attributes of the models tested are summarized in Appendix A.3.
Large Language Models
To gauge the general performance of Large Language Models on bike-bench, we provide an LLM interface with prompts, text descriptions of design parameters and objectives, and a data loader, as described in Appendix A.3.5. We benchmarked OpenAI’s o4-mini-2025-04-16. At the time of benchmarking, May 10, 2025, this was the highest performing model on ArtificalAnalysis’s LLM intelligence index, a blended evaluation suite of seven individual evaluations [55, 56, 57, 58, 59, 60]. To limit cost and benchmarking time, we used the standard inference effort setting, despite ‘high effort’ being used for leaderboard scores. To comply with the model’s context window, we exposed it to 10 example conditions (text, rider geometry, and use case), 10 example designs, and the associated 10 sets of 25 evaluation scores resulting from the condition-design pairs. The model was then provided 10 new conditions (or one in the unconditional setting) and asked to generate a csv file with 10 new designs. During our initial experimentation, it was found that the model tended to duplicate designs. Thus, our prompts include an explicit instruction that each generated design must be unique. The model was queried repeatedly and independently to populate each of 10k + 10k designs used for conditional and unconditional evaluation.
@@
Bike-bench, like many tabular data problems, features ordinal, boolean, and categorical features. Although categorical features can be one-hot encoded into booleans and booleans can be treated as continuous, such simplifications can have significant drawbacks. Training deep learning models directly on categorical data generally requires special modifications, particularly in categorical data generation settings where gradients must be propagated through this discretization. Several generative modeling formulations which learn to generate mixed continuous and categorical data have been proposed over the year. We benchmark two popular models, CTGAN and TVAE, proposed in  [41]. These models are unconditional, do not consider design performance, and are benchmarked principally for their distribution-learning capabilities.
@@
In generative modeling problems related to engineering design, design optimality and diversity are commonly valued in addition to distributional similarity. One mechanism to encourage the generation of high-performing, diverse, and realistic designs it to augment a model’s classic statistical similarity objective with an auxiliary loss factoring in the optimality and diversity of a generated data batch. This involves calling the evaluation functions during training, much like the loop of an optimization algorithm. Since this loss serves to steer the generative model toward more optimal and constraint-satisfying solutions, this procedure effectively constitutes an augmentation-based optimization process. This optimization-augmentation has been demonstrated alongside GAN models using a Determinantal Point Process (DPP)-based auxiliary loss [61]. We benchmark such a performance-augmented GAN and further extend this DPP-based training to a VAE and DDPM. For the DDPM, we test this optimization-augmentation both in training and during inference, as guidance (labeled with ‘-G’). All optimization-augmented benchmarks are labeled with the ‘OA-’ prefix in our benchmarks. More details on our implementation are included in the Appendix.
@@
We evaluate both gradient-based and heuristic optimization. We benchmark plain aggregation-based gradient descent [62] and gradient-based Exact Pareto Optimal (EPA) search [63], both implemented in the LibMoon optimization library [64]. All gradient-based algorithms are applied to one-hot encoded data with constraints enforced using a 1000x weighted penalization function. We also benchmark a mixed-variable adaptation of the Non-dominated Sorting Genetic Algorithm II (NSGA-II) [65], a staple evolutionary optimization algorithm, as implemented in pymoo [66]. Whereas the gradient-based algorithms benefit from the efficiency of gradient information, the mixed-variable NSGA-II learns directly on the mixed-variable data without obfuscation caused by one-hot encoding or continuous relaxation of constraints. Optimization algorithms are only benchmarked on the unconditional test case. All optimization algorithms are run 10 times, one for each of the 10 test conditions. Rather than forcing a final design set of 1k per condition, the final population (featuring less than 1k designs in each algorithm) is taken as the final design set.
@@ 7 Benchmarking Results
In this section, we present and analyze the benchmarking results. Both the unconditional and conditional generation results are compiled in Table 1, with averages over model classes presented in Table 2. All models are evaluated on validity (constraint satisfaction rate), optimality (hypervolume metric), and similarity (maximum mean discrepancy).
Table 1: Full Benchmark Results for unconditional and conditional generation. Models are separated by class (LLM / Tabular-GM / OA-GM / Optimizater), with class-averaged results appearing in Table 2. To provide a grounded reference point for validity and optimality scores, we include a row, dataset, italicized, which benchmarks the trivial ‘generative’ procedure of random selection from the training dataset. High validity and optimality scores are best, while low similarity scores are best. The best non-dataset scores in every metric are bolded.
@@
Validity ($\uparrow$) Optimality ($\uparrow$) Similarity ($\downarrow$) Validity ($\uparrow$) Optimality ($\uparrow$) Similarity ($\downarrow$)
dataset 2.70% 0.2428 0.0006 2.75% 0.4151 0.0007
o4-mini 0.28% 0.0115 0.5026 0.63% 0.2615 0.3460
TVAE 0.05% 0.0334 0.0400 0.00% 0.0000 0.0365
CTGAN 0.58% 0.1496 0.0252 0.87% 0.4090 0.0324
OA-VAE 80.55% 0.2712 0.1081 81.65% 0.4357 0.0693
OA-GAN 24.16% 0.3228 0.1066 22.34% 0.4152 0.1791
OA-DDPM 54.24% 0.2750 0.0740 40.46% 0.3801 0.0590
OA-DDPM-G 40.57% 0.2957 0.0261 40.30% 0.4120 0.0253
Grad-Agg 98.55% 0.3648 0.3824 - - -
EPO 92.00% 0.2625 0.2286 - - -
NSGA-II 100.00% 0.4293 0.5051 - - -
Table 2: Benchmark results averaged by model class
Unconditional Generation Conditional Generation
Val. ($\uparrow$) Opt. ($\uparrow$) Sim. ($\downarrow$) Val. ($\uparrow$) Opt. ($\uparrow$) Sim. ($\downarrow$)
Large Language Model 0.28% 0.0115 0.5026 0.63% 0.2615 0.3460
Tabular Generative Models 0.32% 0.0915 0.0326 0.44% 0.2045 0.0345
Optimization-Augmented Models 49.99% 0.2912 0.0787 46.18% 0.4108 0.0832
Optimization Algorithms 96.85% 0.3522 0.3720 - - -
@@ 7.1 Discussion
The relatively low validity rate of the dataset baseline $(2.7/2.75\%)$ may initially seem surprising. Structural safety factor is a key driver of this low validity: $96\%$ and $84\%$ of dataset designs fail the planar and eccentric safety factor constraints, respectively. This systematic under-engineering is one of the main sources of bias in the BIKED dataset, arising from the fact that many designers never adjust tube thickness parameters, since they are not visually prominent in the BikeCAD user interface [40]. This presents an interesting test for generative models, which, to satisfy structural constraints, generally have to deviate from the norms of the dataset to systematically thicken certain tubes. This exemplifies the principal challenge of Bike-Bench: To simultaneously succeed in validity, optimality, and similarity, models must subtly and tactfully deviate from the distribution of the dataset to achieve design goals.
In general, performance across classes of models is much more variable than performance within classes. We discuss the performance of each class of models one by one. Before doing so, we note a few general comments. First, conditional and unconditional scores should not necessarily be compared. For example, because hypervolume increases with set size given sufficient diversity, we expect significantly higher optimality scores in conditional generation due to the 10x larger set size. We generally expect validity and similarity scores to be relatively similar in unconditional versus conditional generation. Second, for models with very low validity, hypervolume is calculated over very few designs, thus may be somewhat sensitive.
@@
The rich history of bicycle design has extensive textual documentation and undoubtedly appears in common LLM training datasets. Consequently, LLMs can theoretically leverage contextual knowledge of the design parameters and the evaluation criteria to gain an edge over other classes of models. However, they must overcome key disadvantages in their lack of specialization for tabular data, as well as limited context windows, which may prevent them from observing the full dataset. OpenAI’s o4-mini, the lone LLM tested, was a clear underperformer across all metrics, scoring the worst in distributional similarity in all unconditional generation metrics and in similarity during conditional generation. Notably, it did outperform tabular generative models in both validity and optimality during conditional generation, suggesting that its contextual knowledge or exposure to evaluation criteria gave it a slight edge over pure distribution-learning models. o4-mini’s performance indicates, unsurprisingly, that LLMs have room for growth in constrained engineering design problems and tabular data domains.
@@
The tabular generative models trained are unconditional and agnostic to design constraints and objectives. Consequently, they displayed some of the lowest validity and optimality scores, but achieved consistently strong similarity scores. Notably, they underperformed the random dataset sampling baseline in every metric, highlighting the importance of performance-aware training, compared to pure performance-agnostic distribution learning.
@@
Optimization-Augmented Generative models achieved the most balanced set of scores. They demonstrated strong validity and optimality, second only to direct optimization in unconditional tests. Their similarity scores consistently placed them ahead of optimization algorithms and the LLM, but typically behind tabular generative models. The Optimization-Augmented VAE (OA-VAE) achieved high validity scores at the expense of similarity, achieving worst- or second-worst-in-class similarity scores. OA-GAN was generally underwhelming across all metrics. Finally, the guided diffusion model (OA-DDPM-G) was particularly outstanding in its distributional similarity performance, while still maintaining validity and optimality scores commensurate with most other optimization-augmented models. Guided diffusion outperformed unguided conditional diffusion in nearly all metrics.
@@
Optimization algorithms cannot be easily applied in the conditional generation case, thus are limited in their application. This is because they assume evaluation functions are deterministic based solely on the optimization’s design parameters. Since many of Bike-Bench’s evaluation functions depend on conditional information, this conditioning must be static for optimization to succeed. As the only class of non-data-driven methods, optimization is naturally a weak performer in similarity. Noting these limitations, optimization is the clear winner in both validity and optimality for unconditional design generation. The performance of optimization algorithms indicate considerable room for growth for generative models in validity and optimality.

### v2

@@ 5 Benchmarking Metrics and Procedure
The design objectives detailed above yield a set of 10 design objectives and 40 design constraints, which are explained in detail in the Appendix B.4. However, to compare the performance of different generative models, we desire a concise set of summary metrics to evaluate sample sets. BikeBench’s metrics and general benchmarking procedure are introduced in Figure 3, and described in detail in this section.
Figure 3: High-level overview of the benchmarking procedure.
@@ 5.1 Metrics
Though there are countless possibilities for generative model evaluation metrics [24], we select four principal metrics: design quality, constraint violation, similarity to data, and design diversity, which we describe below. Unlike other benchmarks with singular objectives, models benchmarked on BikeBench should be compared across all four scoring metrics. We also evaluate a few auxiliary bonus objectives, as discussed in Appendix B.3.1.
Design Quality (Hypervolume Metric):
High-quality design tools should generate high-quality designs. To quantify design quality, we calculate the hypervolume metric over any designs that simultaneously satisfy all constraints. The hypervolume metric is a staple of multi-objective optimization literature which calculates the overall multi-objective optimality of a set of designs. Because hypervolume directly benefits from constraint satisfaction and diversity, hypervolume can be seen as the best single metric to indicate design utility. More details on the hypervolume metric and its calculation are included in the Appendix B.3.1.
Constraint Violation:
Effective design synthesis tools can reliably navigate design constraints. We measure the mean number of constraints that the model violates per design.
Similarity to Data (Maximum Mean Discrepancy):
Basing new designs on existing ones can be an effective method to minimize risk and maximize feasibility. To capture a model’s ability to match the manifold of existing designs, we evaluate distributional similarity to a subset of 759 designs withheld from the model during training. We select Maximum Mean Discrepancy, a common kernel-based statistical discrepancy measure used to compare sets of samples. Statistical similarity serves to enforce two desirable attributes in generative models for design [24]. The first is ‘design realism’—what does it mean to be a bike and not something entirely different, say, a wheelbarrow? The second desirable attribute is ‘design space coverage,’ encouraging models to learn more than just a small, niche subspace of the overall breadth of bicycle designs, expanding their utility.
Design Diversity (DPP Diversity):
Tools that create a diverse set of design solutions offer the most selection to human users. To capture the diversity of generated designs, we measure a Determinantal Point Process (DPP)-based diversity score, as described in [24]. More details on DPP diversity are included in Appendix B.3.2.
@@ 5.2 Benchmarking Procedure
BikeBench’s benchmarking procedure evaluates 100 designs for each of 100 fixed conditional cases (text prompt, rider anthropometric dimensions, use case). Distributional similarity, constraint satisfaction rate over all 40 constraints, and hypervolume metric over all 10 evaluation objectives are evaluated for each set of 100 bikes, and the mean over the 100 conditions is reported as the final summary score for each metric.
To avoid bias, models are trained only on the 3036 training samples from the original 3798 human-designed bicycles, and not on synthetically generated designs (the synthetic dataset is used exclusively for evaluation of aesthetics scores). Models are allowed to train on either a mixed-datatype or continuous version of the data. Designs generated by models trained on continuous data are mapped back to a mixed-datatype representation prior to evaluation—booleans are decided using rounding, while one-hot categorical data is decided using argmax.
5.3 Benchmark Settings
BikeBench is configurable to provide benchmarking scenarios that accommodate a variety of real-world application cases. This configurability is described below:
Conditional Generation:
In the conditional generation setting, the test-set conditions can only be accessed after all evaluator calls are complete, forcing the method to have conditional generation capabilities (zero-shot generation on new conditions).
Gradient-Free Evaluation:
Although all of BikeBench’s evaluators offer gradients, gradients can be turned off to benchmark gradient-free methods.
Masked Constraints:
To assess models’ ability to infer constraints implicitly from data, we offer a benchmarking case where all but six of the 32 geometric validity constraints are masked to the model. The six retained are the constraints that are violated by at least 5% of the dataset, where the dataset wouldn’t be a good baseline for inferring them.
Evaluation Limits:
Since design evaluation can be costly, design synthesis methods typically try to minimize the number of evaluation calls. BikeBench tallies the number of design evaluations used by each model. Note that for unconditional methods like optimization, evaluations will be summed across all test conditions. To standardize methods, we provide a few max-evaluation brackets: 0, 1K, 1M, 1B, and unlimited.
We establish the “standard” BikeBench settings as follows: Models can call at most 1M gradient-supporting evaluations and will receive completely unmasked scores. However, our leaderboard will support easy filtering of methods and a few preset alternative configurations so that users can easily compare and report scores on non-standard tasks. We request that users who make comparative performance claims on non-standard benchmarks clearly report the BikeBench settings under which they are making these claims. Our benchmark also includes standardized scorecards that showcase key model scores, objective distributions, and constraint satisfaction rates, as shown in Figure 4.
Figure 4: Example scorecard. The scorecard shows key model performance metrics at the top, objective score distributions in the middle, and constraint violation rates at the bottom.
@@ 6 Models Benchmarked
Our principal goal in the benchmarking of baselines is to span a variety of types of design generation procedures ranging from LLMs to tabular generative models to optimization algorithms. Because we do not focus on extensively comparing methods within specific classes, we limit our benchmarking to just one or a few methods per class. The various features and attributes of the models tested are summarized in Appendix B.3.
Baselines:
Our benchmarking starts with two simple baselines. The first is pure random sampling from the dataset, which is completely agnostic to design quality. The second is rejection sampling from the dataset based on evaluated validity scores.
Large Language Models:
To gauge the general performance of Large Language Models on BikeBench, we provide an LLM interface with prompts, text descriptions of design parameters and objectives, and a data loader, as described in Appendix B.3.3. We benchmarked OpenAI’s GPT-5-high. At the time of benchmarking (October 20, 2025) this was the highest performing model on ArtificialAnalysis’s LLM intelligence index, a blended evaluation suite of seven individual evaluations [58, 59, 60, 61, 62, 63]. To comply with the model’s context window, we exposed it to examples of 25 valid designs and 25 invalid designs, their associated conditions, and their corresponding evaluation scores resulting from the condition–design pairs. The model was then provided one new condition and asked to generate a CSV file with the 100 solution designs. This was repeated for the 100 conditions. In the conditional case, the examples provided are for the same condition that the model is generating for.
@@
BikeBench, like many tabular data problems, contains ordinal, boolean, and categorical features. Training deep learning models directly on categorical data generally requires special modifications, particularly in categorical data generation settings where gradients must be propagated through this discretization. Several generative modeling formulations have been proposed to generate mixed continuous and categorical data. We benchmark two popular models, CTGAN and TVAE, proposed in [44]. These models do not consider design performance, and are benchmarked principally for their distribution-learning capabilities. Since these models do not utilize the design evaluation budget during training, the budget is spent on rejection sampling of generated designs.
@@
Performance-augmented training is a popular approach to train generative AI models to synthesize high-quality designs. One method involves calling the design evaluation function during the training process, effectively realizing an optimization-augmented generative model. To balance a tendency toward mode collapse, a diversity-based auxiliary loss can encourage diverse generation of high-performing designs [64]. This technique has subsequently been extended to handle multiple objectives and constraints [52]. We benchmark this optimization-augmented training formulation using a GAN, VAE, and DDPM as base models. We also test a DDPM variant where the auxiliary objective is instead applied during guidance (labeled with -G). All optimization-augmented benchmarks are labeled with the OA- prefix in our benchmarks. More details on our implementation are included in Appendix B.3.2.
@@
We benchmark both gradient-based and heuristic optimization. We test plain aggregation-based gradient descent [65], implemented in the LibMoon optimization library [66]. Gradient descent is applied to one-hot encoded data with constraints enforced using a 1000x weighted penalization function. We also benchmark a mixed-variable adaptation of the Non-dominated Sorting Genetic Algorithm II (NSGA-II) [67], a staple evolutionary optimization algorithm, as implemented in pymoo [68]. Whereas the gradient-based algorithms benefit from the efficiency of gradient information, the mixed-variable NSGA-II operates directly on the mixed-variable data without obfuscation caused by one-hot encoding or continuous relaxation of constraints. All optimization algorithms are run once for each of the 100 test conditions, splitting the evaluation budget. Optimization is initialized per-condition using valid designs from the dataset.
@@ 7 Benchmarking Results
In this section, we present and analyze the benchmarking results using the aforementioned “standard” evaluation procedure: unmasked constraints, gradients, and a 1M evaluation budget. Results in both unconditional and conditional generation are compiled in Table 1. All models are evaluated on design quality, constraint violation, similarity to dataset, and diversity. Detailed scorecards for all tested models are included in Appendix A.1.
Table 1: Design quality, constraint violation, similarity to dataset, and design diversity scores for unconditional and conditional benchmarking cases. Models are separated by class (LLM / Tabular-GM / OA-GM / Optimizer). The best non-baseline scores in each metric are bolded. Models are benchmarked on unmasked constraints with an evaluation budget of 1M.
@@
Qual. ($\uparrow$) Viol. ($\downarrow$) Sim. ($\downarrow$) Div. ($\downarrow$) Qual. ($\uparrow$) Viol. ($\downarrow$) Sim. ($\downarrow$) Div. ($\downarrow$)
Dataset 0.0035 2.699 0.005 12.72 0.0035 2.699 0.005 12.72
Data subset 0.0151 0.000 0.081 12.28 0.0151 0.112 0.079 12.43
GPT-5-high 0.0086 0.448 0.363 14.08 0.0049 1.288 0.631 14.70
CTGAN 0.0087 0.000 0.147 7.10 0.0148 0.318 0.082 11.93
TVAE 0.0160 0.000 0.163 13.55 0.0161 0.053 0.153 13.55
OA-GAN 0.0137 0.000 0.308 14.44 0.0109 2.579 0.079 14.04
OA-VAE 0.0181 0.000 0.301 14.96 0.0186 0.001 0.348 14.97
OA-DDPM 0.0168 0.000 0.077 12.55 0.0165 0.042 0.084 12.82
OA-DDPM-G 0.0195 0.000 0.037 13.03 - - - -
Grad. Descent 0.0164 0.092 0.266 12.04 - - - -
NSGA-II 0.0102 0.000 0.622 11.88 - - - -
@@ 7.1 Discussion
Baselines:
The relatively high constraint violation of the dataset baseline (avg. 2.7 constraints per design) may initially seem surprising. The structural safety factor is a key driver of this low validity: $95\%$ and $80\%$ of dataset designs fail the planar and eccentric safety factor constraints, respectively. This systematic under-engineering is one of the main sources of bias in the BIKED dataset, arising from the fact that many designers never adjust tube thickness parameters, since they are not visually prominent in the BikeCAD user interface [43]. This presents an interesting test for generative models, which, to satisfy structural constraints, generally have to deviate from the norms of the dataset to systematically thicken certain tubes. This exemplifies the principal challenge of BikeBench: To simultaneously succeed across metrics, models must subtly but strategically deviate from the distribution of the dataset to achieve design goals. Although sampling only valid designs from the dataset is a fairly strong baseline, strong models should be able to outperform it.
@@
The rich history of bicycle design has extensive textual documentation and almost certainly appears in common LLM training datasets. Consequently, LLMs can theoretically leverage contextual knowledge of the design parameters and the evaluation criteria to gain an edge over other classes of models. However, they must overcome key disadvantages in their lack of specialization for tabular data, as well as limited context windows, which may prevent them from observing the full dataset. OpenAI’s GPT-5-high, the lone LLM tested, was a clear underperformer across all metrics, ranking in the bottom three methods for each objective. This indicates, unsurprisingly, that LLMs have substantial room for growth in constrained engineering design problems and tabular-data domains.
@@
The tabular generative models trained were unconditional and agnostic to design constraints and objectives. The evaluations were only invoked to sample subsets of generated designs. Across the board, their design quality, constraint violation, and similarity to data were generally equal to or worse than the simple baseline of dataset subsampling. However, CTGAN was a notable frontrunner in diversity score. These benchmarking results highlight the importance of performance-aware training, compared to subsampling of pure performance-agnostic generated samples.
@@
Optimization-Augmented Generative Models achieved the best quality, constraint violation, and similarity scores in both unconditional and conditional generation. While the guided DDPM was especially dominant in the unconditional setting, the standard DDPM was a strong performer across the board. The VAE was a strong performer in quality but its similarity and diversity scores were much worse than the DDPMs. Finally, the GAN underperformed its optimization-augmented counterparts in most metrics. In general, the OA-GMs achieved strong balanced scores with standout performance from optimization-augmented DDPM models.
@@
Classic optimization algorithms were relatively unimpressive performers on BikeBench. However, their diversity scores were notably strong, indicating a capability to maintain a wider variety of high-quality solutions compared to competing methods. Interestingly, the diversity of AI-generated design sets is often one of their selling points over classically-optimized design sets. BikeBench’s results indicate that excessive fixation on design quality can cause AI-generated designs to be less diverse than classically-optimized designs. Gradient-based optimization achieved reasonably strong design quality scores with decent similarity. NSGA-II struggled in both design quality and dataset similarity. The performance of these classic optimization algorithms highlights a major shortcoming: Even when amortizing evaluation budget over just 100 designs, generative models outperform optimization in classic optimization metrics. This suggests considerable room for growth, perhaps using approaches that would amortize learning across conditional cases.